Can someone help me solve this by elemination? What is the solution?
x + 5y = 24
-x+ 5y = 6

Found 2 solutions by rwm, sudhanshu_kmr:
Answer by rwm(914)

add them
10y=30
y=3
x+5*3=24
x=9

Answer by sudhanshu_kmr(1152)


add both equations.
10 y = 30
=> y = 30/10 = 3
put this value of first equation,
x + 15 = 24
=> x = 24-15 =9
so, x= 9, y = 3


Everyone is welcome to get free online help: sudhanshu.cochin@gmail.com